# Quantitative 2D propagation of smallness and control for 1D heat equations with power growth potentials

### v1

@@ 4.2. Propagation of smallness: nondivergence form
When $\varepsilon C_{1}<1$, we must have $|\omega_{a}|\leq|\omega|/2$. Indeed, otherwise (4.26) would hold with $\omega$ replaced by $\omega_{a}$, leading to
@@
Let $\gamma_{n}\in(0,\frac{1}{2}),n\in\mathbb{Z}$ and $V\in\mathcal{C}(\mathbb{R})$ be a potential as in Assumption A1. Let $\Omega$ be a measurable subset of  $\mathbb{R}$, $\omega_{n}:=\Omega\cap I_{n}$ such that $|\omega_{n}|=\gamma_{n}|I_{n}|$ for every $n\in\mathbb{Z}$. Then for any real-valued $H^{2}_{\mathrm{loc}}$ solution $\Phi$ of (4.17) in $\mathbb{R}^{2}$, we have
@@
with $d:=d(c_{2},\beta_{2})>0$ depending only on $c_{2}$ and $\beta_{2}$. Notice that
@@
Hence the function $\Phi$ is a $H^{2}_{\mathrm{loc}}$ solution for (4.17) with $\Phi(x,0)=\phi(x)$ on $\mathbb{R}$. Applying Corollary 4.6 to $\Phi$, we obtain
@@
where we have used the inequality $a_{n}^{\frac{\alpha_{n}-2}{2}}\leq 1$. For the second term on the right-hand side of (4.58), we have the following lemma:
@@
where the second inequality uses Lemma 4.1. By the orthogonality of $\phi_{k}$s, we obtain
@@
(4.72) $\sum_{n\in\mathcal{J}}\|\chi_{n}\Phi\|^{2}_{H^{2}(\mathbb{R}^{2})}=$ $\underbrace{\sum_{n\in\mathcal{J}}\|\chi_{n}\Phi\|^{2}_{L^{2}(\mathbb{R}^{2})}}_{\rm S_{1}}+\underbrace{\sum_{n\in\mathcal{J}}\|D_{x}(\chi_{n}\Phi)\|^{2}_{L^{\infty}(\mathbb{R}^{2})}}_{\rm S_{2}}+\underbrace{\sum_{n\in\mathcal{J}}\|D_{y}(\chi_{n}\Phi)\|^{2}_{L^{\infty}(\mathbb{R}^{2})}}_{\rm S_{3}}$
$+$ $\underbrace{\sum_{n\in\mathcal{J}}\|D_{x}^{2}(\chi\Phi)\|^{2}_{L^{2}(\mathbb{R}^{2})}}_{\rm S_{4}}+\underbrace{\sum_{n\in\mathcal{J}}\|D_{y}^{2}(\chi_{n}\Phi)\|^{2}_{L^{2}(\mathbb{R}^{2})}}_{\rm S_{5}}$
@@
We need to check that each one of the five terms is bounded by $e^{5\lambda}$ multiplied by a positive constant. Estimate of $\rm S_{1}$. This term is the simplest one, just extending the integral from $\sum_{n\in\mathcal{J}}\int_{I_{3,n}}$ to $\int_{\mathbb{R}}$ and using the orthogonality of eigenfunctions, we obtain
@@
Let $\gamma_{n}\in(0,\frac{1}{2})$,$n\in\mathbb{Z}$ and $V\in\mathcal{C}(\mathbb{R})$ be the potential as in Assumption A1. Let $\Omega$ be a measurable subset of $\mathbb{R}$ such that $|\omega_{n}|=\gamma_{n}|I_{n}|$ for every $n\in\mathbb{Z}$. Then for any $\varepsilon>0$, there exists a positive number $R_{0}=R(\varepsilon,V)>0$ and a positive integer $n_{0}=n_{0}(R)>0$, such that for any $0<R<R_{0}$ and any real-valued $H^{2}_{\mathrm{loc}}$ solution $\Phi$ of (4.17) in $\mathbb{R}^{2}$, we have
@@
(4.95) $a_{n}=\frac{|x_{n}|^{\frac{\beta_{2}}{2}}}{\left(R\log\log\langle x_{n}\rangle-R^{-1}\right)^{\frac{1}{2}}}$
@@
(4.96) $\widetilde{V}(x)\leq c_{2}\frac{1}{a_{n}^{2}}\left\langle\frac{x}{a_{n}}\right\rangle^{\beta_{2}}\leq C^{\prime}\left(R\log\log\langle x_{n}\rangle-R^{-1}\right)$
for all $x\in a_{n}I_{3,n}$ and for all $|n|$ large enough, where $C^{\prime}=C^{\prime}(c_{2},\beta_{2})$ depends only on $c_{2}$ and $\beta_{2}$. Here we have also used the relation in Lemma 4.8 to approximately identify ${x}/{a_{n}}\in I_{3,n}$ with $x_{n}$. Using Proposition 4.5 to (4.36), we obtain
@@
(A.9) $\|v\|_{\small H^{1}(P_{1}(N))}\leq\delta^{-\alpha}\!\exp\left(C\left(1+\mathcal{G}(V_{1},V_{2},9\sqrt{d}N)\right)\right)\|v\|^{\alpha}_{\small H^{1}(P_{2}(N))}\left\|\frac{\partial v}{\partial y_{d+1}}\right\|^{1-\alpha}_{\small L^{2}(D_{\delta}(N))}\!\!,$
@@ Acknowledgements
The author warmly thanks Philippe Jaming for his constant guidance. He is also particularly grateful to Ming Wang for his suggestive advice and idea about using Zhu-Zhuge’s spectral inequality. The author also thanks Pei Su and Yuzhe Zhu for helping me understand deeply the techniques used in the article. It was supported by China Scholarship Council grant CSC202206410007.

### v2

@@ 4.2. Propagation of smallness: nondivergence form
When $\varepsilon C_{1}<1$, we must have $|\omega_{a}|\leq|\omega|/2$. Indeed, otherwise (4.30) would hold with $\widetilde{\omega}$ replaced by $\omega_{a}$, leading to
@@
Let $\gamma_{n}\in(0,\frac{1}{2}),n\in\mathbb{Z}$ and $V\in L^{\infty}_{\mathrm{loc}}(\mathbb{R})$ be a potential as in Assumption A1. Let $\Omega$ be a measurable subset of  $\mathbb{R}$, $\omega_{n}:=\Omega\cap I_{n}$ such that $|\omega_{n}|=\gamma_{n}|I_{n}|$ for every $n\in\mathbb{Z}$. Then for any real-valued $H^{2}_{\mathrm{loc}}$ solution $\Phi$ of (4.21) in $\mathbb{R}^{2}$, we have
@@
with $d:=d(c_{2},\beta_{2})>0$ depending only on $c_{2}$ and $\beta_{2}$.
Notice that
@@
Hence the function $\Phi$ is a $H^{2}_{\mathrm{loc}}$ solution for (4.21) with $\Phi(x,0)=\phi(x)$ on $\mathbb{R}$.
Applying Corollary 4.5 to $\Phi$, we obtain
@@
where we have used the inequality $a_{n}^{\frac{\alpha_{n}-2}{2}}\leq 1$.
For the second term on the right-hand side of (4.62), we have the following lemma:
@@
where the second inequality uses Lemma 4.1. By the orthogonality of $\phi_{k}$, we obtain
@@
(4.76) $\sum_{n\in\mathcal{J}}\|\chi_{n}\Phi\|^{2}_{H^{2}(\mathbb{R}^{2})}=$ $\underbrace{\sum_{n\in\mathcal{J}}\|\chi_{n}\Phi\|^{2}_{L^{2}(\mathbb{R}^{2})}}_{\rm S_{1}}+\underbrace{\sum_{n\in\mathcal{J}}\|D_{x}(\chi_{n}\Phi)\|^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(\mathbb{R}^{2})}}}_{\rm S_{2}}+\underbrace{\sum_{n\in\mathcal{J}}\|D_{y}(\chi_{n}\Phi)\|^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(\mathbb{R}^{2})}}}_{\rm S_{3}}$
$+$ $\underbrace{\sum_{n\in\mathcal{J}}\|D_{x}^{2}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\chi_{n}}\Phi)\|^{2}_{L^{2}(\mathbb{R}^{2})}}_{\rm S_{4}}+\underbrace{\sum_{n\in\mathcal{J}}\|D_{y}^{2}(\chi_{n}\Phi)\|^{2}_{L^{2}(\mathbb{R}^{2})}}_{\rm S_{5}}$
@@
We need to check that each one of the five terms is bounded by $e^{5\lambda}$ multiplied by a positive constant.
Estimate of $\rm S_{1}$. This term is the simplest one, just extending the integral from $\sum_{n\in\mathcal{J}}\int_{I_{3,n}}$ to $\int_{\mathbb{R}}$ and using the orthogonality of eigenfunctions, we obtain
@@
Let $\gamma_{n}\in(0,\frac{1}{2})$,$n\in\mathbb{Z}$ and $V\in L^{\infty}_{\mathrm{loc}}(\mathbb{R})$ be the potential as in Assumption A1. Let $\Omega$ be a measurable subset of $\mathbb{R}$ such that $|\omega_{n}|=\gamma_{n}|I_{n}|$ for every $n\in\mathbb{Z}$. Then for any $\varepsilon>0$, there exists a positive number $R_{0}=R(\varepsilon,V)>0$ and a positive integer $n_{0}=n_{0}(R)>0$, such that for any $0<R<R_{0}$ and any real-valued $H^{2}_{\mathrm{loc}}$ solution $\Phi$ of (4.21) in $\mathbb{R}^{2}$, we have
@@
(4.99) $a_{n}=\frac{|x_{n}|^{\frac{\beta_{2}}{2}}}{R\log\log\langle x_{n}\rangle-R^{-1}}$
@@
(4.100) $\widetilde{V}(x)\leq c_{2}\frac{1}{a_{n}^{2}}\left\langle\frac{x}{a_{n}}\right\rangle^{\beta_{2}}\leq{C^{\prime}}^{2}\left(R\log\log\langle x_{n}\rangle-R^{-1}\right)^{2}$
for all $x\in a_{n}I_{3,n}$ and for all $|n|$ large enough, where $C^{\prime}=C^{\prime}(c_{2},\beta_{2})$ depends only on $c_{2}$ and $\beta_{2}$. Here we have also used the relation in Lemma 4.7 to approximately identify ${x}/{a_{n}}\in I_{3,n}$ with $x_{n}$.
Using Proposition 4.4 to (4.40), we obtain
@@
(A.9) $\|v\|_{H^{1}(P_{1}(N))}\leq\delta^{-\alpha}\!\exp\left(C\left(1+\mathcal{G}(V_{1},V_{2},9\sqrt{d}N)\right)\right)\|v\|^{\alpha}_{H^{1}(P_{2}(N))}\left\|\frac{\partial v}{\partial y_{d+1}}\right\|^{1-\alpha}_{L^{2}(D_{\delta}(N))}\!\!,$
@@ Acknowledgements
The author warmly thanks Philippe Jaming for his constant guidance. He is also particularly grateful to Ming Wang for his suggestive advice and idea about using Zhu-Zhuge’s spectral inequality. The author also thanks Pei Su, Chenmin Sun and Yuzhe Zhu for helping me understand deeply the techniques used in the article. It was supported by China Scholarship Council grant CSC202206410007.